OpenStudy (anonymous):

1. x – 3 > –9 2. –2x + 1 ≤ –11 3. 10 < –3x + 1 4. 2(x + 5) > 8x – 8 5. –2(x – 3) ≥ 5 – (x + 3) Im having problems especially with 4 and 5. Can someone help me and show me the steps?

OpenStudy (mathstudent55):

4. Start by distributing the 2 through the parentheses. Can you do that?

OpenStudy (anonymous):

Um. Would i add the 2 on each sides?

OpenStudy (mathstudent55):

Forget this problem now. You need to review the distributive property. The distributive property is a(b + c) = ab + ac Example: 2(5 + 6) = 2*5 + 2*6 = 10 + 12 = 22

OpenStudy (mathstudent55):

To use the distributive property, you multiply the number outside the parentheses by each number inside the parentheses.

OpenStudy (mathstudent55):

Now let's get back to your problem.

OpenStudy (anonymous):

I dont have time to review. I have to get 5 other assignments done by today. And this is part of one. I just dont get how to solve inequalities.

OpenStudy (mathstudent55):

4. Start by distributing the 2 through the parentheses. 2(x + 5) > 8x – 8 Using the distributive property on the left side, you get 2x + 2*5 > 8x - 8 2x + 10 > 8x - 8

OpenStudy (mathstudent55):

Ok, now subtract 8x from both sides.

OpenStudy (mathstudent55):

-6x + 10 > -8 Now subtract 10 from both sides: -6x > -18 Now you need to divide both sides by -6 to get x. In an inequiality, when you multiply or divide both sides by a negative number, the inequality sign changes direction: x < 3 Final answer.

OpenStudy (anonymous):

Okay. I think i get it now. So would 1. be x > -6

OpenStudy (anonymous):

And would 2. be x ≥ 6?

OpenStudy (mathstudent55):

1 and 2 are correct

OpenStudy (mathstudent55):

How about 3. First add 3x to both sides to bring the x term to the left side.

OpenStudy (mathstudent55):

Then subtract 10 from both sides.

OpenStudy (anonymous):

-9 > x?

OpenStudy (mathstudent55):

3. First add 3x to both sides to bring the x term to the left side. 10 < –3x + Subtract 10 from both sides: 3x + 10 < 1 3x < -9 Now divide both sides by 3